OpenStudy (anonymous):

x^4-81 divided by x-3

OpenStudy (anonymous):

using synthetic division

OpenStudy (anonymous):

I am not personally smart enough to figure it out, but this site really helped me with this sort of thing: ww.algebrahelp.com/calculators/equation/ Just put your question in there. Today it wasn't working, but I've used it in the past.

OpenStudy (anonymous):

okay Thank you!! (:

OpenStudy (anonymous):

Let me know if it doesn't work :) and your welcome :)

OpenStudy (anonymous):

yeah i don't think its working :(

OpenStudy (anonymous):

Ok, I know the site wasn't working today, but if you don't need it right away, you can try again later. Sorry :(

OpenStudy (anonymous):

well you an factorize the equation, because you know that 9^2= 81, :\[x ^{4} - 81 = (x ^{2}-9)*(x ^{2}+9)\] it is a property of factorization-->\[x ^{2}- a ^{2}=(x-a)(x+a)\].. follow this you can factorize also \[x ^{2}-9 = (x+3)(x-3)\] and then you can to elimine (x-3) in the denominator and the numerator...

OpenStudy (anonymous):

at te end, you have \[\frac{ (x-3)*(x+3)*(x ^{2}+9) }{x-3 }= (x+3)*(x ^{2}+9)\]